OpenStudy (anonymous):

find the critical points: s(t)= (t-1)^4 (t+5)^3

OpenStudy (anonymous):

first we find the derivative.. i know that.. then...?

OpenStudy (anonymous):

then equate it to 0

OpenStudy (jennychan12):

set it = 0

OpenStudy (anonymous):

but wait, im confused to find the derivative for this one..

OpenStudy (anonymous):

s'(t) =0

OpenStudy (jennychan12):

chain rule/product rule

OpenStudy (anonymous):

its 4(t-1)^3 3(t+5)^2

OpenStudy (anonymous):

Yup

OpenStudy (anonymous):

what's next after it? i'm a little lost after it

OpenStudy (anonymous):

4(t-1)^3 3(t+5)^2 = 0

OpenStudy (anonymous):

Now Find t

OpenStudy (anonymous):

what about chain rule?

OpenStudy (jennychan12):

doesnt apply here because derviative of t is just 1

OpenStudy (anonymous):

if ur function is of the type (t^2 + 2)^4 then s'(t) = 4 (t^2+2) * 2t

OpenStudy (anonymous):

sorry s'(t) = 4 (t^2+2)^3 * 2t

OpenStudy (anonymous):

oh right.

OpenStudy (anonymous):

so if the derivative of t is 1 then how do i solve t?

OpenStudy (anonymous):

sorry imm not sure what to do after ... 4(t-1)^3 3(t+5)^2